### v1

@@ IV-C Ablation Analysis
Quantum vs. Classical: The top-3 performing models are all quantum-based, with QTCNN achieving 72% improvement over the best classical baseline (Transformer). This suggests that quantum feature encoding provides complementary representational capacity for cross-sectional ranking tasks.
@@ V-A Summary
V-B Limitations and Future Work
@@
Hardware noise characterization. All experiments in this study assume ideal, noiseless quantum simulation. Near-term quantum devices are subject to various noise sources including gate errors, decoherence, and measurement noise, which may significantly degrade model performance. Characterizing the noise resilience of QTCNN through simulation with realistic noise models (e.g., depolarizing channels, amplitude damping) and eventual deployment on quantum hardware remains an important direction for future investigation.
@@
[Uncaptioned image] Chi-Sheng Chen received a bachelor degree in Electrophysics from National Chiao Tung University and a master’s degree in Biomedical Electronics and Bioinformatics from National Taiwan University. He is currently working at a financial AI startup and also serves as an AI project researcher at Harvard Medical School and Beth Israel Deaconess Medical Center. His research expertise includes time-series analysis, computer vision, brain-computer interfaces (BCI), electroencephalography (EEG), multimodal contrastive learning, and quantum machine learning. Chen has led the development of foundational models for EEG-based image generation, aiming to improve psychiatric treatment and the discovery of biomarkers for drug development. He has designed hybrid quantum-classical architectures for EEG encoding and pioneered quantum deep learning methods for contrastive learning between EEG and visual data. His research also focuses on medical speech processing and the integration of multimodal large language models (LLMs) for clinical and financial application.
[Uncaptioned image] Xinyu Zhang received the B.E. and the M.S. degree from the Information and Automatic College, Tianjin University of Science and Technology, Tianjin, China, in 2019 and 2022, separately, where he is currently pursuing the Ph.D. degree in intelligent system engineering with Luddy School of Informatics, Computing, and Engineering, Indiana University Bloomington, Bloomington, IN, USA. His main research includes quantum neural network, deep learning, advanced signal processing, and other disciplines such as information science and large language models (LLMs).
[Uncaptioned image] Rong Fu received the B.S. degree in Communication Engineering from Jiangsu Ocean University, Jiangsu, China, in 2018, and the M.S. degree in Electronic Information Engineering from the College of Information and Automation, Tianjin University of Science and Technology, Tianjin, China, in 2023. She is currently pursuing the Ph.D. degree in Intelligent Systems Engineering with the Luddy School of Informatics, Computing, and Engineering, Indiana University Bloomington, Bloomington, IN, USA. Her research interests include signal processing and deep learning.
[Uncaptioned image] Qiuzhe Xie received a bachelor’s degree in Electrical Engineering from Hefei University of Technology and a master’s degree in Electronics Engineering from National Cheng Kung University, followed by a Ph.D. in Electronics Engineering from National Taiwan University. His research expertise encompasses biosensors, advanced semiconductor processes, nanomaterials, and microfabrication techniques. He has spearheaded the development of nano-structural electrode arrays with ultra-thin dielectric stacking and leveraged quantum computing algorithms to enhance sensing precision. Furthermore, he has engineered CRISPR-based electrochemical sensors and developed innovative water purification systems for semiconductor manufacturing. Extending his work to practical application, he also focuses on the commercialization of medical devices, having validated business models for cardiosensor technologies through the NSF program.
[Uncaptioned image] Fan Zhang received the B.S. degree in Statistics from Beijing Normal University, Beijing, China, in 2016, the M.S. degree in statistics from National Cheng Kung University, Taiwan, in 2018, and the Ph.D. degree in statistics from Arizona State University, Tempe, AZ, USA, in 2024. She is currently an Assistant Professor with the Department of Mathematics, Boise State University, Boise, ID, USA. Her research interests include design and analysis of experiments, variable selection, uncertainty quantification, Bayesian modeling and inference, and quantum-enhanced machine learning.
[…]

### v2

@@ IV-C4 Architecture-Level Comparisons
Quantum vs. Classical: The top-4 performing models (QTCNN, QNN, QTCNN-NoEnt, TConv-MLP) all include the temporal encoder, while the top-2 (QTCNN, QNN) both use entangling quantum circuits, reinforcing the complementary value of both components.
IV-D Statistical Significance
Given the limited 22-day OOS period, establishing statistical significance requires careful analysis. We employ two complementary methods.
IV-D1 Paired Bootstrap Test
For each model pair (QTCNN vs. competitor), we compute the daily long-short return difference $\delta_{t}=r^{\text{QTCNN}}_{\text{LS},t}-r^{\text{comp}}_{\text{LS},t}$ for $t=1,\ldots,22$. We then resample $\{\delta_{t}\}$ with replacement B=10, [^0] B=10{,}000 times and compute the fraction of bootstrap replicates where the resampled mean $\leq 0$. The resulting $p$-values are reported in Table VI. QTCNN significantly outperforms the Transformer ($p=0.002$), QLSTM ($p=0.003$), and TConv-MLP ($p=0.019$) at the 5% level. The comparison with QTCNN-NoEnt ($p=0.038$) is also significant, confirming that entanglement provides a statistically reliable improvement. The comparison with QNN ($p=0.048$) is marginally significant, while the comparison with QCNN ($p=0.182$) is not, reflecting QCNN’s high variance ($\pm 0.298$).
IV-D2 Ledoit-Wolf Robust Sharpe Ratio Test
We additionally apply the robust Sharpe ratio test of Ledoit and Wolf [ledoit2008robust], which accounts for non-normality and serial correlation in returns. Under this test, the null hypothesis $H_{0}:\text{SR}_{\text{QTCNN}}=\text{SR}_{\text{Transformer}}$ is rejected with $p=0.027$ (HAC kernel with Parzen weights, bandwidth selected via Andrews’ method). For the QTCNN vs. TConv-MLP comparison, the Ledoit-Wolf $p$-value is $0.041$, confirming significance under a more conservative testing framework.
IV-D3 Power Analysis
[…]
IV-E Qubit Scalability Study
To evaluate the persistence of the quantum advantage across different problem sizes, we conduct a scalability study by varying the qubit count $n_{q}\in\{4,6,8,10,12\}$. We compare the top-performing QTCNN against the pure QNN baseline, keeping all other hyperparameters fixed. The PCA projection dimension is adjusted to match $n_{q}$ for each experiment.
Table VII summarizes the results. We observe that the Sharpe ratio for QTCNN improves monotonically from $n_{q}=4$ (0.347) to $n_{q}=10$ (0.551), where it reaches a plateau. At $n_{q}=12$, we observe a slight decline in performance (0.529), which is consistent with the onset of barren plateaus and increased training difficulty in larger Hilbert spaces. The marginal gain from $n_{q}=8$ to $n_{q}=10$ is only 2.4%, suggesting that the 8-qubit regime already captures the essential cross-sectional structure of the returns. Notably, QTCNN consistently outperforms QNN at every qubit count, confirming that the advantage of the temporal encoder is orthogonal to the quantum circuit capacity. While classical simulation time grows exponentially ($O(2^{n_{q}})$), the gate count on quantum hardware remains polynomial, demonstrating a viable path toward larger-scale financial models.
TABLE VII: Qubit Scalability Analysis for QTCNN and QNN.
$n_{q}$ Hilbert Dim Q-Params Gates QTCNN Sharpe QNN Sharpe Time/iter (s)
[…]
IV-F Mechanism of Quantum Enhancement
The superior performance of QTCNN over classical baselines can be attributed to several quantum-mechanical mechanisms:
Hilbert space expressibility: The angle embedding $\Phi:\mathbb{R}^{n_{q}}\to\mathbb{C}^{2^{n_{q}}}$ maps $n_{q}$ classical features into an exponentially larger space [sim2019expressibility]. For $n_{q}=8$, the quantum circuit operates in a 256-dimensional complex Hilbert space, whereas a classical linear model with 8 features spans only an 8-dimensional real space. The variational layers learn to exploit this exponentially richer representation to identify subtle alpha signals.
Entanglement and feature interaction: Each CNOT gate in the ring topology creates quantum entanglement between qubits encoding different financial features. For example, momentum encoded on qubit $i$ becomes entangled with volatility on qubit $j$. This enables the circuit to represent joint multi-feature interactions that would require explicit higher-order polynomial terms or significantly deeper architectures in classical models. To quantify this, we compute the Meyer-Wallach entanglement measure $Q_{MW}$ [meyer2002global] for QTCNN circuits. At random initialization, $Q_{MW}=0.412\pm 0.089$, which increases to $0.687\pm 0.054$ after training, indicating that the model learns to exploit quantum correlations for prediction.
Quantum interference as implicit feature selection: The variational $R_{Y}/R_{Z}$ rotations create constructive interference, amplifying probability amplitudes for feature combinations predictive of returns, and destructive interference, suppressing noisy or irrelevant patterns. This process is analogous to learned attention weights in Transformers but operates within the exponentially larger Hilbert space.
Implicit regularization from quantum geometry: The parameter-shared QTCNN uses only 18 quantum parameters to access a 256-dimensional Hilbert space. This extreme compression acts as a strong geometric prior, forcing the model to explore a low-dimensional manifold of the full Hilbert space. This naturally prevents overfitting, explaining why QTCNN outperforms both the under-regularized QCNN (42 parameters, higher variance) and classical models that require explicit regularization.
IV-G Gradient Variance and Trainability Analysis
A critical concern for variational quantum algorithms is the barren plateau phenomenon [mcclean2018barren], whereby the variance of parameter gradients vanishes exponentially with qubit count, rendering optimization infeasible. We empirically characterize the trainability of QTCNN by measuring the gradient variance across different circuit widths and comparing against architectures without parameter sharing.
IV-G1 Measurement Protocol
For each architecture and qubit count $n_{q}\in\{4,6,8,10,12\}$, we initialize 200 independent random parameter configurations drawn uniformly from $0,2\pi)$. For each configuration, we compute the gradient $\partial\mathcal{L}/\partial\theta_{1}$ with respect to the first variational parameter using the parameter-shift rule (Equation 52) on a fixed batch of 128 training samples. We then report $\text{Var}_{\bm{\theta}}[\partial\mathcal{L}/\partial\theta_{1}]$ across the 200 initializations.
IV-G2 Results
Table VIII presents the measured gradient variance for three quantum architectures: QTCNN (parameter-shared convolution), QCNN (independent parameters per pair), and QNN (BasicEntanglerLayers).
TABLE VIII: Gradient variance $\text{Var}[\partial\mathcal{L}/\partial\theta_{1}]$ across 200 random initializations. QTCNN exhibits polynomial decay; QCNN and QNN exhibit exponential decay consistent with barren plateaus.
$n_{q}$ QTCNN QCNN QNN
4 $8.23\times 10^{-2}$ $7.56\times 10^{-2}$ $6.89\times 10^{-2}$
6 $3.71\times 10^{-2}$ $1.98\times 10^{-2}$ $1.62\times 10^{-2}$
8 $1.84\times 10^{-2}$ $5.1\times 10^{-3}$ $3.8\times 10^{-3}$
10 $9.7\times 10^{-3}$ $1.2\times 10^{-3}$ $8\times 10^{-4}$
12 $5.9\times 10^{-3}$ $3\times 10^{-4}$ $2\times 10^{-4}$
IV-G3 Analysis
Three observations are noteworthy:
  1. 1.
[…]
  2. 2.
    Parameter sharing is the key factor. The primary difference between QTCNN and QCNN is intra-layer parameter sharing. By constraining all qubit pairs within a layer to use identical parameters, QTCNN restricts the circuit to a low-dimensional manifold that avoids the concentration-of-measure effects responsible for barren plateaus. This is consistent with theoretical results showing that shallow, structured circuits can evade exponential gradient vanishing [cerezo2021variational].
  3. 3.
    Practical training threshold. Gradient variance below $\sim 10^{-3}$ typically renders stochastic gradient descent ineffective for standard batch sizes. For QTCNN, this threshold is not reached until $n_{q}>20$ (extrapolating the polynomial fit), suggesting that QTCNN remains trainable up to $\sim$20 qubits without requiring specialized initialization or layerwise training strategies. In contrast, QCNN and QNN reach this threshold at $n_{q}\approx 10$$12$, consistent with the performance plateau observed in Table VII.
This trainability advantage complements QTCNN’s representational advantage identified in Section IV-D and partially explains the consistent performance gap between QTCNN and other quantum architectures across all qubit counts (Table VII).
IV-H Noise Resilience and Hardware Feasibility
To assess the prospects for practical deployment on NISQ hardware, we conduct a systematic noise resilience study by simulating QTCNN and QNN under two physically motivated noise channels.
IV-H1 Simulation Methodology
All noisy simulations are performed using PennyLane’s default.mixed backend, which evolves the full $2^{n_{q}}\times 2^{n_{q}}$ density matrix $\rho$ rather than a pure state vector. This backend supports the insertion of arbitrary quantum channels within the circuit. Specifically, we insert noise channels after every quantum gate (both single-qubit rotations and two-qubit CNOT gates) using PennyLane’s built-in channel operations:
  • Depolarizing channel (qml.DepolarizingChannel): After each gate acting on qubit $j$, we apply
    $\mathcal{E}_{\text{dep}}(\rho)=(1-p)\,\rho+\frac{p}{3}\bigl(X_{j}\rho X_{j}+Y_{j}\rho Y_{j}+Z_{j}\rho Z_{j}\bigr),$ (59)
    where $p$ is the single-qubit depolarizing probability. For two-qubit CNOT gates, the channel is applied independently to each of the two participating qubits at the same error rate $p$.
  • Amplitude damping channel (qml.AmplitudeDamping): After each gate on qubit $j$, we apply the channel with Kraus operators
    $K_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\quad K_{1}=\begin{bmatrix}0&\sqrt{\gamma}\\ 0&0\end{bmatrix}$, (60)
    where $\gamma$ is the damping parameter. This models energy relaxation ($T_{1}$ decay) in superconducting qubits.
For each noise configuration, we retrain the full model from scratch using the same hyperparameters (Table IV) and report the mean $\pm$ standard deviation of the Sharpe ratio across 5 independent runs.
IV-H2 Choice of Error Rates
The error rates $p\in\{10^{-3},\;5\times 10^{-3},\;10^{-2},\;2\times 10^{-2}\}$ are chosen to span the range of gate fidelities reported by current superconducting quantum processors:
[…]
  • $p=5\times 10^{-3}$: typical of median-quality qubits on current 100+ qubit devices, where not all qubit pairs achieve best-case fidelity.
  • $p=10^{-2}$: represents the threshold below which variational algorithms are generally expected to produce meaningful results without error correction [cerezo2021variational].
  • $p=2\times 10^{-2}$: a pessimistic scenario probing the graceful degradation regime.
[…]
IV-H3 Results and Analysis
Table IX presents the out-of-sample Sharpe ratio under each noise configuration. Several observations are noteworthy:
  1. 1.
    Noise tolerance threshold: QTCNN maintains a Sharpe ratio exceeding the best classical baseline (Transformer: 0.313) up to a depolarizing error rate of $p\approx 0.01$. This indicates that near-term devices with two-qubit gate fidelities $\geq 99\%$ are sufficient for practical deployment.
  2. 2.
    QTCNN vs. QNN robustness: QTCNN exhibits consistently greater noise resilience than QNN across all error rates, with the gap widening as $p$ increases. We attribute this to QTCNN’s shallower effective circuit depth after pooling: at $n_{q}=8$, the final measurement involves only 1 qubit (after 3 pooling stages), whereas QNN measures all 8 qubits, accumulating more noise.
  3. 3.
[…]
  4. 4.
    Practical device regime: At $p=10^{-3}$, which matches state-of-the-art device calibration data, QTCNN’s Sharpe ratio degrades by only 3.2% (depolarizing) or 1.7% (amplitude damping), suggesting that the ideal-simulation results are a reliable proxy for near-term hardware performance. Error mitigation techniques such as zero-noise extrapolation (ZNE) [temme2017error] and probabilistic error cancellation (PEC) [li2017efficient] could further close this gap without requiring full quantum error correction.
TABLE IX: Out-of-sample Sharpe ratio of QTCNN and QNN under simulated gate noise ($n_{q}=8$, default.mixed backend). $\Delta$ denotes the relative change from the ideal (noiseless) QTCNN baseline.
Noise Model Error Rate $p$ QTCNN Sharpe QNN Sharpe $\Delta$ QTCNN (%)
[…]
@@ V-A Summary
V-B Dequantization Considerations
At $n_{q}=8$, the quantum state can be exactly classically simulated as a 256-dimensional state vector; accordingly, we do not claim a computational quantum speedup in this simulation regime. Instead, we interpret the observed advantage as representational: the angle embedding followed by entangling layers creates an implicit feature map $\Phi:\mathbb{R}^{8}\to\mathbb{C}^{256}$ whose inductive bias is naturally aligned with the structure of cross-sectional equity returns. A classical neural network with comparable capacity (e.g., 256 hidden units) would require substantially more trainable parameters and training data to learn a similar inductive bias from scratch.
Recent dequantization literature, such as Tang (2019) [tang2019quantum], has shown that certain quantum machine learning algorithms can be dequantized when the input data has low-rank structure. However, our angle embedding does not fall into the amplitude-encoding framework that most dequantization results target [schuld2021effect]. The critical question for scalability is whether the performance gains observed at $n_{q}=8-10$ persist to larger qubit counts. At $n_{q}\geq 20$, classical state vector simulation becomes intractable ($2^{20}>10^{6}$ amplitudes). If the monotonic improvement shown in Table VII continues, the quantum advantage would become both representational and computational.
V-C Prospects for NISQ and Fault-Tolerant Execution
[…]
From our noise analysis (Table IX), QTCNN requires a two-qubit gate fidelity $\geq 99\%$ (error rate $\leq 0.01$) to exceed classical baselines. Current state-of-the-art devices from IBM and Google already achieve fidelities in the range of 99.5–99.9%. Scaling to a production system processing $\sim$2000 stocks with $n_{q}=16$ qubits would require $\sim$106 gates per sample. With hardware throughput reaching $10^{4}$ circuits per second, daily prediction for the entire universe would take approximately 200 seconds, which is competitive with classical GPU inference.
In the fault-tolerant regime, under surface code error correction with a physical error rate of $10^{-3}$, each logical qubit requires approximately 1000 physical qubits. A 16-logical-qubit QTCNN would thus need $\sim$16,000 physical qubits, which is within the long-term roadmaps of major quantum hardware providers [gambetta2020ibm]. For near-term deployment, error mitigation techniques such as zero-noise extrapolation (ZNE) [temme2017error] and probabilistic error cancellation (PEC) [li2017efficient] can further extend the noise tolerance, making QTCNN viable on today’s hardware.
V-D Limitations and Future Work
@@
Extended hardware noise characterization. While Section IV-E provides an initial noise resilience study under depolarizing and amplitude damping channels, a full characterization of QTCNN under device-specific noise profiles—including crosstalk, readout errors, and non-Markovian decoherence—remains an important direction for future work. Eventual deployment on physical quantum hardware will be necessary to validate the simulated noise results and to assess the practical effectiveness of error mitigation techniques under realistic operating conditions.